OpenStudy (anonymous):

write an equation in slope-intercept form of a line containing the points( 1/2,6) and (-1/2,8) a)y=-2x+7 b) y=-x+5 c)y=-x+7 d)y=-2x+5

OpenStudy (anonymous):

a

OpenStudy (angela210793):

slope=(y2-y1)/(x2-x1)=-1

OpenStudy (angela210793):

How??

OpenStudy (angela210793):

-1/2-1/2=-2/2=-1

OpenStudy (anonymous):

y-6=-2(x-1/2) y=-2x+1+6=-2x+7

OpenStudy (angela210793):

i think iy is c

OpenStudy (anonymous):

m=6-8/(1/2+1/2)=-2

OpenStudy (angela210793):

huh? isn't m=(y2-y1)/(x2-x1)

OpenStudy (anonymous):

yes

OpenStudy (angela210793):

then??? How did u get -2?

OpenStudy (anonymous):

what did u get?

OpenStudy (angela210793):

-1

OpenStudy (anonymous):

(1/2,6), (-1/2,8) are the points

OpenStudy (anonymous):

6-8/(1/2+1/2)=-2.....no?

OpenStudy (angela210793):

yea so 8-6=2 and -1/2-1/2=-1 and ur right

OpenStudy (angela210793):

OOPS!!!

OpenStudy (anonymous):

:)